OpenStudy (anonymous):

Jenny, Demi, and Drew buy an 18-in. long gourmet sandwich and take it back to their apartment. Since they have different appetites, Jenny cuts the sandwich so that Demi gets half of what Jenny gets and Drew gets three-fourths of what Jenny gets. Find the length of each sandwich?

OpenStudy (mertsj):

x=jenny's 1/2x=Demi's 3/4x=Drews 9/4x=18 x=8 Jenny gets 8 inches, demi gets 4 inches and drew gets 6 inches